OpenStudy (anonymous):

The area of a square is (4x^2-12x+9) sq.cm.The width of a rectangle is five cm greater than a side of the square while the length of the rectangle is twice the side of the square.What is the area, in sq.cm. of the rectangle? please show me the steps. A.16x^2-12x-14 B.8x^2+10x-12 C.4x^2-8x-16 D.4x^2+2x-2 E.None of these

OpenStudy (blockcolder):

One side of the square is given by sqrt(4x^2-12x+9)=2x-3 cm. If L and W are the length and width of the rectangle, then L=2x+2 cm. and W=4x-6 cm. Now what is the area of the rectangle?

OpenStudy (anonymous):

@blockcolder @sammy12 block is correct!! but block u can make it more easier!

OpenStudy (blockcolder):

@Rohangrr You mean like this: LW=(2x+2)(4x-6)=4(x+1)(2x-3) ??

OpenStudy (anonymous):

Is it 8x^2-4x-12 The answer is E .Is it Right!

OpenStudy (blockcolder):

Yes. Yes it is. :D

OpenStudy (anonymous):

Thank you so much

OpenStudy (blockcolder):

No problem. :D